Section 5. Problem 13 (25 pts.) For the below beam the loads are: 10 (k), 3 (k-ft.). 9 (k) and 5 (k). The reactions for the bSaved H Ak Ck B k/ft. Dk A B с D Ε 6 ft. 6 ft. 4 ft. 4 ft.
Answer #1

Date lok 9K 1X ski 3Kft B e D Efy IRB RE Jo+36 +9+5 RBARE oft 44- 6st 457 RBTRE = 60 28.8571 K (10.8574 - 1.857e E 13-1436 lV (B) right Ver to for Shear force bet ne (o cross-section at distance & from & xx). Vex = -10 -3 0 + RB -9 VCA) = -lok V (B